# EvolveNav: Self-Improving Embodied Reasoning for LLM-Based Vision-Language Navigation

### v2

@@ 3.2 Model Architecture
We build the LLM-based navigation agent that can simultaneously produce the navigational chain-of-thought reasoning and action prediction, modifying from a recent LLM-based VLN work, NaviLLM zheng2024towards . Concretely, the navigation agent consists of a scene encoder $F_{v}$, an LLM backbone $F_{\mathrm{LLM}}$, and an action prediction head $F_{\mathrm{action}}$. At each timestep $t$, the agent receives the navigation instruction $I$, panoramic observation $O_{t}$, and navigation history features $H_{t}=\{h_{0},...,h_{t-1}\}$. The scene encoder transforms $N$ navigable panoramic views $\{O_{t,n}\}_{n=1}^{N}$ into visual representations $\{V_{t,n}\}_{n=1}^{N}$:
@@
The navigation prompt $P$ is then constructed by integrating the tokenized instruction, the visual representations $\{V_{t,n}\}_{n=1}^{N}$, and navigation history features $H_{t}$. As shown in Figure 2, special tokens <hist> and <cand> are introduced as placeholder tokens, where we insert the features $H_{t}$ and $\{V_{t,n}\}_{n=1}^{N}$, respectively. In contrast to NaviLLM zheng2024towards that directly maps the navigational inputs to action decision, in our EvolveNav, we construct the following output hint in the prompt $P$ to guide the navigation agent to generate both the action decision and explicit navigational reasoning: “-Action Decision: <cls>. -Navigational Reasoning: ”. The <cls> token is also a special token for facilitating subsequent action predictions. The navigation prompt $P$ is fed into the LLM backbone $F_{\mathrm{LLM}}$ to obtain the feature ${f}^{cls}_{t}$ of the <cls> token and the chain-of-thought (CoT) reasoning $\mathrm{CoT}$:
@@
3.3 Self-Improving Embodied Reasoning
@@ 3.3.1 Stage 1: Formalized CoT Supervised Fine-Tuning
Formalized CoT Labels Collection. When facing a given human instruction, the navigation agent usually needs to sequentially reason about the direction or the landmark it should move to in its current visual observation to reach the target position. Therefore, in our EvolveNav, we train the LLM-based VLN model to generate the chain-of-thought (CoT) reasoning about the landmark with the corresponding direction at different navigation timesteps, like the following format: “I should go to an observation with [landmark] to the [direction] of me”.
To encourage the navigation agent to choose the ground-truth action $a^{*}_{t}$ (paired observation is denoted as $O_{t}^{*}$) at different timesteps $t$ through the guidance of CoT reasoning during training, we obtain the corresponding landmarks $L$ and direction $D$ of $O_{t}^{*}$ to construct the formalized CoT labels, which is described as follows. Denote the observation $O_{t}^{*}$ as $O_{t}^{*}=\{B_{t},A_{t}=\{\psi_{t},\theta_{t}\}\}$, where $B_{t}$ is the RGB image of $O_{t}^{*}$, $A_{t}$ represent the direction information containing heading $\psi_{t}$ and elevation $\theta_{t}$. For the image $B_{t}$, we first employ a powerful image captioning model huang2023tag2text $F_{\mathrm{cap}}$ to obtain object and scene context $C_{t}$:
@@
Then, we leverage the NLP tool Spacy spacy to extract the landmarks list $L$ from $C_{t}$. In contrast to directly using object recognition models which may detect multiple redundant objects, extracting landmarks from the image captions can better retain salient landmarks. As a result, the generated CoT reasoning of the navigation agent can effectively help it locate important landmarks mentioned in the human instruction, since humans also tend to focus on salient landmarks when giving navigation instructions. We follow lin2024navcot to map the direction information $A_{t}$ of the observation $O_{t}^{*}$ to textual represented direction $D$. With the landmarks list $L$ and direction $D$, we construct CoT labels $\mathrm{\mathrm{CoT}^{*}}$ by filling the following label template: I should go to an observation with [$L$] to the [$D$] of me. Through using our formalized CoT labels for training, the generation of redundant reasoning information can be effectively alleviated to promote the reasoning speed.
@@
where $\lambda$ represents the loss balance factor. We follow zheng2024towards to calculate the action prediction training objective $\mathcal{L}_{\mathrm{action}}$.
@@ 3.3.2 Stage 2: Self-Reflective Post-Training
Although CoT supervised fine-tuning can explicitly guide the navigation agent to produce navigational reasoning for assisting the action decision, due to the uncertainty and complexity of the navigation task, using fixed labels may lead to overfitting to training CoT label distributions and therefore harm the generalization to unseen scenarios. Moreover, the inherent noise in the image captioning model huang2023tag2text for landmark detection may also limit the accuracy of formalized CoT labels collected in Stage 1. Therefore, after the Stage 1 training of Formalized CoT Supervised Fine-Tuning, we introduce Self-Reflective Post-Training to further encourage the navigation agent to learn correct reasoning patterns for improving generalization.
@@
Self-Reflective Auxiliary Task. To further make the model aware of correct and wrong reasoning, which can help the model better learn correct reasoning patterns, we additionally introduce a self-reflective auxiliary task, where we ask the model to discriminate which reasoning output from the given reasoning is right. Specifically, we collect positive and negative reasoning samples $R^{+}$ and $R^{-}$ during training for conducting the self-reflective auxiliary task. We utilize the above mentioned CoT label $\tilde{\mathrm{CoT}}_{t}^{*}$ as the positive reasoning sample $R^{+}$. To obtain the negative reasoning sample $R^{-}$, we randomly select the candidate observation (action) $O_{t,j}$ ($1<j<N$, $N$ is the number of navigable views) different from the ground-truth one. Then we extract the landmark and direction for $O_{t,j}$ to fill in the CoT label template (see Sec. 3.3.1). As shown in Figure 2, we construct the self-reflective task prompt $P_{\mathrm{sr}}$ as “Choose the correct one from the given two navigational reasoning outputs. Output 1: [$R^{1}$]. Output 2: [$R^{2}$]. Selection: ”, where we randomly insert the positive reasoning sample $R^{+}$ and negative reasoning sample $R^{-}$ to the positions of $R^{1}$ and $R^{2}$. We collect the ground-truth output $R_{\mathrm{sr}}^{*}$ with the form like “Output 2.”. We also utilize the auto-regressive training objective like Eq. 5 to calculate the loss $\mathcal{L}_{\mathrm{sr}}$ based on ($P_{\mathrm{sr}}$, $R_{\mathrm{sr}}^{*}$) pairs for the self-reflective auxiliary task.
@@ 4.1 Experimental Setup
Datasets. We test our EvolveNav on two popular VLN benchmarks, i.e., R2R anderson2018vision and CVDN thomason2019vision . R2R is built on 90 real-world indoor simulation environments containing 7,189 trajectories, each corresponding to three fine-grained instructions. CVDN contains 2,050 human-human navigation dialogs and over 7k trajectories in 83 MatterPort houses.
Evaluation Metrics. We utilize the following standard metrics for evaluation: 1) Trajectory Length (TL): the average length of the agent’s navigated path, 2) Navigation Error (NE): the average distance between the agent’s destination and the goal viewpoint, 3) Success Rate (SR): the ratio of success, where the agent stops within three meters of the target point, 4) Success rate weighted by Path Length (SPLanderson2018vision : success rate normalized by the ratio between the length of the shortest path and the predicted path, 5) Oracle Success Rate (OSR): the ratio of containing a viewpoint along the path where the target position is visible, 6) Goal Progress (GP), the progress in meters towards the goal.
Table 1: Comparison with state-of-the-art methods on the R2R dataset. * denotes training with R2R in-domain data. IL means the imitation learning setting.
Method Val Unseen Test Unseen
@@
PREVALENT hao2020towards 10.19 4.71 58 53 10.51 5.30 54 51
HOP qiao2022hop 12.27 3.80 64 57 12.68 3.83 64 59
HAMT Chen2021HistoryAM 11.46 2.29 66 61 12.27 3.93 65 60
VLN-BERT hong2021vln 12.01 3.93 63 57 12.35 4.09 63 57
DUET Chen2022ThinkGA 13.94 3.31 72 60 14.73 3.65 69 59
Meta-Explore Hwang2023MetaExploreEH 13.09 3.22 72 62 14.25 3.57 71 61
VLN-SIG li2023improving - - 72 62 - - 72 60
VLN-PETL qiao2023vln 11.52 3.53 65 60 12.30 4.10 63 58
NavGPT2 zhou2024navgpt 13.25 3.18 71 60 - - - -
@@
NavGPT zhou2023navgpt 11.45 6.46 34 29 - - - -
DiscussNav long2023discuss 9.69 5.32 43 40 - - - -
MapGPT chen2024mapgpt - 5.63 34 29 - - - -
NavCoT lin2024navcot 9.95 6.26 40 37 - - - -
NaviLLM zheng2024towards 12.81 3.51 67 59 13.21 3.71 68 60
NaviLLM* (IL) 9.99 6.04 46.90 43.78 10.03 6.12 46 43
EvolveNav (IL) 9.79 5.52 51.15 48.27 9.94 5.92 47 45
NaviLLM* 13.43 3.27 70.11 60.25 13.68 3.37 70 61
EvolveNav 12.07 3.15 71.17 63.48 12.06 3.22 71 63
Implementation Details. We fine-tune the LLM with full-parameter and LoRA settings for Stage 1 and 2, respectively. To simplify the implementation, we do not introduce the pretraining phase like zheng2024towards for both our EvolveNav and the compared baseline (denoted as NaviLLM* in Table 1 &  2). The training for Stage 1 is conducted on 8 Nvidia A100 GPUs and the training for Stage 2 is performed on 4 Nvidia A100 GPUs. Empirically, we set the loss coefficients $\lambda$, $\lambda_{1}$, and $\lambda_{2}$ as 1, 1, and 0.2, respectively. The hyperparameters such as the learning rate, optimizer, and the batch size are kept the same as zheng2024towards .More implementation details are presented in the supplementary material.
@@ 4.2 Comparison with Existing Methods
Table 1 shows the comparison among different approaches on R2R. We compare EvolveNav with the baseline method NaviLLM zheng2024towards using the R2R in-domain training data (denoted as NaviLLM*) under two settings, i.e., imitation learning (IL) and Dagger training zheng2024towards ; Chen2022ThinkGA . The comparison shows the consistent improvement of our approach over NaviLLM* on both Val Unseen and Test Unseen under different training settings. For example, the performance gain in SPL of EvolveNav on Val Unseen is $\sim$4.5and $\sim$3.2under IL and dagger training settings compared to the baseline, respectively. Note that our EvolveNav achieves the state-of-the-art performance on R2R among the approaches with LLM-based backbone. These results demonstrate the effectiveness of our self-improving embodied reasoning framework in fulfilling generalizable navigational reasoning. We do not consider the comparison with models augmented by new environments (e.g., ScaleVLN wang2023scaling ) for fairness.
Table 2: Comparison with SoTA methods on CVDN. We utilize the Goal Progress (GP) (m) as the evaluation metric. * denotes training with CVDN in-domain data.
Method Val-Unseen Test
Cross-Modal Backbone:
Seq2Seq thomason2019vision 2.10 2.35
PREVALENT hao2020towards 3.15 2.44
HOP qiao2022hop 4.41 3.31
MT-RCM wang2020environment 4.36 -
MT-RCM+Env wang2020environment 4.65 3.91
HAMT Chen2021HistoryAM 5.13 5.58
VLN-SIG li2023improving 5.52 5.83
VLN-PETL qiao2023vln 5.69 6.13
LLM-based Backbone:
NaviLLM zheng2024towards 6.16 7.90
NaviLLM* 5.53 6.80
EvolveNav (ours) 6.21 7.07
In Table 2, we compare our EvolveNav with different VLN approaches on CVDN. As shown in Table 2, EvolveNav outperforms the baseline NaviLLM* zheng2024towards on both Val Unseen and Test set on CVDN, demonstrating good generalization ability of our proposed self-improving embodied reasoning paradigm to different kinds of VLN tasks. For example, EvolveNav brings $\sim$0.7 performance gain over NaviLLM* in GP on Val Unseen. Moreover, Table 2 shows the superiority of EvolveNav over NaviLLM zheng2024towards , which conducts multi-task training on multiple VLN benchmarks to learn cross-task navigation knowledge, demonstrating the importance of explicit navigational reasoning learning compared to direct input-output mapping paradigms. Such results also reveal the effectiveness of our constructed CoT labels, which provide landmark and direction information for enabling the agent to learn useful navigational reasoning for decisions. The results on both R2R and CVDN benchmarks demonstrate the effectiveness of our EvolveNav in fulfilling generalizable navigational reasoning to improve navigation decision accuracy.

### v3

@@ 3.2 Model Architecture
We build the LLM-based navigation agent that can simultaneously produce the navigational chain-of-thought reasoning and action prediction, modifying from a recent LLM-based VLN work, NaviLLM [45]. The navigation agent consists of a scene encoder $F_{v}$, an LLM backbone $F_{\mathrm{LLM}}$, and an action prediction head $F_{\mathrm{action}}$. At each timestep $t$, the agent receives the navigation instruction $I$, panoramic observation $O_{t}$, and navigation history features $H_{t}=\{h_{0},...,h_{t-1}\}$. The scene encoder transforms $N$ navigable panoramic views $\{O_{t,n}\}_{n=1}^{N}$ into visual representations $\{V_{t,n}\}_{n=1}^{N}$:
@@
The navigation prompt $P$ is then constructed by integrating the tokenized instruction, the visual representations $\{V_{t,n}\}_{n=1}^{N}$, and navigation history features $H_{t}$. As shown in Figure 2, special tokens <<hist>> and <<cand>> are introduced as placeholder tokens, where we insert the features $H_{t}$ and $\{V_{t,n}\}_{n=1}^{N}$, respectively.
In contrast to NaviLLM [45] that directly maps the navigational inputs to action decision, in EvolveNav, we construct the following output hint in the prompt $P$ to guide the navigation agent to generate both the action decision and explicit navigational reasoning: “-Action Decision: <<cls>>. -Navigational Reasoning: ”. The <<cls>> token is also a special token for facilitating subsequent action predictions. The navigation prompt $P$ is fed into the LLM backbone $F_{\mathrm{LLM}}$ to obtain the feature ${f}^{cls}_{t}$ of the <<cls>> token and the chain-of-thought (CoT) reasoning $\mathrm{CoT}$:
@@
3.3 Self-Improving Embodied Reasoning Framework
@@ 3.3.1 Stage 1: Formalized CoT Supervised Fine-Tuning
Formalized CoT Labels Collection. When facing a given human instruction, the navigation agent usually needs to sequentially reason about the direction or the landmark it should move to in its current visual observation to reach the target position. Therefore, in EvolveNav, we train the LLM-based VLN model to generate the chain-of-thought (CoT) reasoning about the landmark with the corresponding direction at different navigation timesteps, like the following format: “I should go to an observation with [landmark] to the [direction] of me”.
To encourage the navigation agent to choose the ground-truth action $a^{*}_{t}$ (paired observation is denoted as $O_{t}^{*}$) at different timesteps $t$ through the guidance of CoT reasoning during training, we obtain the corresponding landmarks $L$ and direction $D$ of $O_{t}^{*}$ to construct the formalized CoT labels, which is described as follows. Denote the observation $O_{t}^{*}$ as $O_{t}^{*}=\{B_{t},A_{t}=\{\psi_{t},\theta_{t}\}\}$, where $B_{t}$ is the RGB image of $O_{t}^{*}$, $A_{t}$ represent the direction information containing heading $\psi_{t}$ and elevation $\theta_{t}$. For the image $B_{t}$, we first employ a powerful image captioning model [60] $F_{\mathrm{cap}}$ to obtain object and scene context $C_{t}$:
@@
Then, we leverage the NLP tool Spacy [61] to extract the landmarks list $L$ from $C_{t}$. In contrast to directly using object recognition models which may detect multiple redundant objects, extracting landmarks from the image captions can better retain salient landmarks. As a result, the generated CoT reasoning of the navigation agent can effectively help it locate important landmarks mentioned in the human instruction, since humans also tend to focus on salient landmarks when giving navigation instructions. We follow [32] to map the direction information $A_{t}$ of the observation $O_{t}^{*}$ to textual represented direction $D$. With the landmarks list $L$ and direction $D$, we construct CoT labels $\mathrm{\mathrm{CoT}^{*}}$ by filling the following label template: I should go to an observation with [$L$] to the [$D$] of me.
Through extracting the landmark and direction information of the ground-truth observation (action) straightforwardly to construct the CoT labels, we do not explicitly correlate the CoT labels and the navigation instructions, leading to excellent generalization to navigational instructions of various types. Such CoT label construction strategy can effectively alleviate the problem that some action decisions are not explicitly corresponding to the navigation instruction, while instead enabling latent alignment learning of action decision, CoT reasoning, and navigational inputs.
@@
where $\lambda$ represents the loss balance factor. We follow [45] to calculate the action prediction training objective $\mathcal{L}_{\mathrm{action}}$.
Merits of formalized CoT labels. Our design of formalized CoT labels has the following merits compared with free-form CoT labels like those collected in [43] (also see Figure 2): Firstly, as the VLN task needs sequential decision making, generating formalized CoTs can significantly promote the reasoning speed compared to generating free-form ones. Secondly, free-form CoTs created by modern frontier models like GPT-4 [43] may produce irrelevant and redundant reasoning for decision, while formalized CoT can produce concise and task-related reasoning. Thirdly, using formalized CoT labels for training can effectively simplify the training process as well as mitigate the hallucination compared to using free-form ones.
TABLE I: Performance comparison results on R2R under the task-specific training setting. * denote our reimplementation results. IL means the imitation learning setting. The best results for Cross-Modal Backbone and LLM-based Backbone are annotated in blue and bold fonts, respectively.
Method Val Unseen Test Unseen
[…]
Cross-Modal Backbone:
PREVALENT [26] 10.19 4.71 58 53 10.51 5.30 54 51
HOP [62] 12.27 3.80 64 57 12.68 3.83 64 59
HAMT [13] 11.46 2.29 66 61 12.27 3.93 65 60
VLN-BERT [12] 12.01 3.93 63 57 12.35 4.09 63 57
DUET [15] 13.94 3.31 72 60 14.73 3.65 69 59
Meta-Explore [63] 13.09 3.22 72 62 14.25 3.57 71 61
VLN-SIG [64] - - 72 62 - - 72 60
VLN-PETL [65] 11.52 3.53 65 60 12.30 4.10 63 58
NavGPT2 [43] 13.25 3.18 71 60 - - - -
LLM-based Backbone:
NavGPT [24] 11.45 6.46 34 29 - - - -
DiscussNav [25] 9.69 5.32 43 40 - - - -
MapGPT [23] - 5.63 34 29 - - - -
NavCoT [32] 9.95 6.26 40 37 - - - -
NaviLLM* [45] (IL) 9.99 6.04 46.90 43.78 10.03 6.12 46 43
EvolveNav (IL, ours) 9.79 5.52 51.15 48.27 9.94 5.92 47 45
NaviLLM* [45] 13.43 3.27 70.11 60.25 13.68 3.37 70 61
EvolveNav (ours) 12.07 3.15 71.17 63.48 12.06 3.22 71 63
@@ 3.3.2 Stage 2: Self-Reflective Post-Training
Although CoT supervised fine-tuning can explicitly guide the navigation agent to produce navigational reasoning for assisting the action decision, due to the uncertainty and complexity of the navigation task, using fixed labels may lead to overfitting to training CoT label distributions and therefore harm the generalization to unseen scenarios. Moreover, the inherent noise in the image captioning model [60] for landmark detection may also limit the accuracy of formalized CoT labels collected in Stage 1. Therefore, after the Stage 1 training of Formalized CoT Supervised Fine-Tuning, we introduce Self-Reflective Post-Training to further encourage the navigation agent to learn correct reasoning patterns in a self-improving manner for improving generalization.
@@
Self-Reflective Auxiliary Task. To further make the model aware of correct and wrong reasoning, which can help the model better learn correct reasoning patterns, we additionally introduce a self-reflective auxiliary task, where we ask the model to discriminate which reasoning output from the given reasoning is right. Specifically, we collect positive and negative reasoning samples $R^{+}$ and $R^{-}$ during training for conducting the self-reflective auxiliary task. We utilize the above mentioned CoT label $\tilde{\mathrm{CoT}}_{t}^{*}$ as the positive reasoning sample $R^{+}$. To obtain the negative reasoning sample $R^{-}$, we randomly select the candidate observation (action) $O_{t,j}$ ($1<j<N$, NN is the number of navigable views) different from the ground-truth one. Then we extract the landmark and direction for $O_{t,j}$ to fill in the CoT label template (see Sec. 3.3.1). As shown in Figure 2, we construct the self-reflective task prompt $P_{\mathrm{sr}}$ as “Choose the correct one from the given two navigational reasoning outputs. Output 1: $R^{1}$]. Output 2: $R^{2}$]. Selection: ”, where we randomly insert the positive reasoning sample $R^{+}$ and negative reasoning sample $R^{-}$ to the positions of $R^{1}$ and $R^{2}$. We collect the ground-truth output $R_{\mathrm{sr}}^{*}$ with the form like “Output 2.”. We also utilize the auto-regressive training objective like Eq. 5 to calculate the loss $\mathcal{L}_{\mathrm{sr}}$ based on ($P_{\mathrm{sr}}$, $R_{\mathrm{sr}}^{*}$) pairs for the self-reflective auxiliary task.
@@
3.3.3 Training and Inference
Two-stage training. When conducting training in the proposed self-improving embodied reasoning framework, we train the model in Stage 1 to converge, and use it for Stage 2 training. This two-stage training design is to mitigate the negative impact of noisy CoT outputs during Stage 1 under a non-converged situation and ensure the training stability during Stage 2. Specifically, in the early training iterations during Stage 1 of Formalized CoT Supervised Fine-Tuning, the agent is prone to generate CoT reasoning with noisy formats and information (e.g., its output may contain repeatedly notions of “<<s>>” or the output may be very long) while accompanied action decisions may be correct occasionally. Based on the rule of our self-enriched label enhancement strategy during Stage 2, such noisy outputs will be introduced as CoT labels to guide the CoT training and may cause training instability. By firstly training the model in Stage 1 to converge, such issues can be effectively alleviated. This two-stage training manner also simplifies the method design and implementation to promote its practicality during real deployment.
Inference. During inference, the model generates CoT based on the given prompt with the form of “-Action Decision: <<cls>>. -Navigational Reasoning:”. The encoding of the <<cls>> token is extracted to generate the action prediction probability $a_{t}$ through the action prediction head $F_{\mathrm{action}}$ (see Sec. 3.2).
TABLE II: Performance comparison results on CVDN under the task-specific training setting. We utilize the Goal Progress (GP) (m) as the evaluation metric. * denote our reimplementation results. The best results for Cross-Modal Backbone and LLM-based Backbone are annotated in blue and bold fonts, respectively.
Method Val-Unseen Test
Cross-Modal Backbone:
Seq2Seq [34] 2.10 2.35
PREVALENT [26] 3.15 2.44
HOP [62] 4.41 3.31
MT-RCM [66] 4.36 -
MT-RCM+Env [66] 4.65 3.91
HAMT [13] 5.13 5.58
VLN-SIG [64] 5.52 5.83
VLN-PETL [65] 5.69 6.13
LLM-based Backbone:
NaviLLM* [45] 5.53 6.80
EvolveNav (ours) 6.21 7.07
@@ 4.1 Experimental Setup
4.1.1 Datasets
We test EvolveNav on four popular VLN benchmarks, i.e., R2R [33], CVDN [34], REVERIE [35], and SOON [36]. Each benchmark handles distinct challenges posed by VLN. R2R is built on 90 real-world indoor simulation environments containing 7,189 trajectories, each corresponding to three fine-grained instructions. CVDN contains 2,050 human-human navigation dialogs and over 7k trajectories in 83 MatterPort houses. REVERIE replaces the fine-grained instructions in R2R with high-level instructions. SOON constructs thoroughly described instructions to further highlight visual-semantic alignment.
To verify the effectiveness of EvolveNav, we adopt two representative training setting, task-specific training and cross-task training. Task-specific training trains the model on single benchmark like most previous works [13, 15, 32], while cross-task training realizes a generalist navigation model like NaviLLM [45] by training the model using multiple datasets.
4.1.2 Evaluation Metrics
We utilize the following standard metrics for evaluation: 1) Trajectory Length (TL): the average length of the agent’s navigated path, 2) Navigation Error (NE): the average distance between the agent’s destination and the goal viewpoint, 3) Success Rate (SR): the ratio of success, where the agent stops within three meters of the target point, 4) Success rate weighted by Path Length (SPL[33]: success rate normalized by the ratio between the length of the shortest path and the predicted path, 5) Oracle Success Rate (OSR): the ratio of containing a viewpoint along the path where the target position is visible, 6) Goal Progress (GP), the progress in meters towards the goal.
4.1.3 Implementation Details
We utilize NaviLLM [45] as our baseline model. To simplify the implementation, we do not introduce the pretraining phase in [45] for both our EvolveNav and the baseline (denoted as NaviLLM* in Table IV). In our EvolveNav, We fine-tune the LLM with full-parameter and LoRA settings for Stage 1 and 2 training, respectively. The training for Stage 1 is conducted on 8 Nvidia A100 GPUs and the training for Stage 2 is performed on 4 Nvidia A100 GPUs. Empirically, we set the loss coefficients $\lambda$, $\lambda_{1}$, and $\lambda_{2}$ as 1, 1, and 0.2, respectively. During Stage 1 training, we introduce the CoT supervised finetuning loss $\mathcal{L}_{\mathrm{SFT}}$ under a probability of 0.5 to mitigate the overfitting to the pre-constructed CoT labels. The maximum numbers of training steps for Stage 1 and 2 are set as 60000 and 9000 steps, respectively. Training for Stage 1 with full parameter lasts for $\sim$1.5 days with $\sim$73G GPU memory, and training for Stage 2 with LoRA lasts for $\sim$1 day with $\sim$30G GPU memory. The hyperparameters such as the learning rate, optimizer, and the batch size are kept the same as [45].
TABLE III: Performance comparison results on SOON under the task-specific training setting. * denote our reimplementation results. The best results for Cross-Modal Backbone and LLM-based Backbone are annotated in blue and bold fonts, respectively.
Method Val-Unseen
@@ 4.1.3 Implementation Details
GBE [36] 28.54 19.52 13.34
DUET [15] 50.91 36.28 22.58
AZHP [67] 56.19 40.71 26.58
@@
NaviLLM* [45] 43.60 30.34 23.70
EvolveNav (Ours) 49.56 33.40 24.92
TABLE IV: Performance comparison results on REVERIE under the task-specific training setting. * denote our reimplementation results. The best results for Cross-Modal Backbone and LLM-based Backbone are annotated in blue and bold fonts, respectively.
Method Val Unseen
[…]
Cross-Modal Backbone:
Seq2Seq [33] 8.07 4.20 2.84
HOP [62] 36.24 31.78 26.11
HAMT [13] 36.84 32.95 30.20
VLN-BERT [12] 35.02 30.67 24.90
DUET [15] 51.07 46.98 33.73
AZHP [67] 53.65 48.31 36.63
VLN-PETL [65] 37.03 31.81 27.67
LLM-based Backbone:
NaviLLM* [45] 42.68 32.55 25.82
EvolveNav (Ours) 42.40 33.60 28.16
@@ 4.2 Comparison with Existing Methods
Task-specific training. Table IIIIII, and IV present the task-specific training results on R2R [33], CVDN [34], SOON [36], and REVERIE [35], respectively. The results exhibit consistent superiority of EvolveNav over the compared approaches on various VLN benchmarks, demonstrating the effectiveness and excellent generalization ability of the proposed self-improving embodied reasoning paradigm. For example, for the results on R2R in Table I, the performance gain in SPL of EvolveNav on Val Unseen is $\sim$4.5% and $\sim$3.2% under the imitation learning (IL) and dagger [15] training settings compared to the baseline model NaviLLM [45], respectively. For the results on SOON in Table [36], the performance improvements of EvolveNav over the baseline model NaviLLM on OSR, SR, and SPL are $\sim$5.9%, $\sim$3.1%, and $\sim$1.2%, respectively. Note that we do not consider the comparison with models augmented by new environments (e.g., ScaleVLN [41]) for fairness.
TABLE V: Performance comparison results under the cross-task training setting. * denote our reimplementation results. The best results for task-specific training (“Separate Model for Each Task”) and cross-task training (“Unified Model for All Tasks”) are annotated in blue and bold fonts, respectively.
Method REVERIE SOON R2R CVDN
[…]
Separate Model for Each Task:
DUET [15] 46.98 51.07 33.73 36.28 50.91 22.58 72 - 60 -
AZHP [67] 48.31 53.65 36.63 40.71 56.19 26.58 - - - -
VLN-PETL [65] 31.81 37.03 27.67 - - - 65 - 60 5.69
NaviLLM* [45] 32.55 42.68 25.82 30.34 43.60 23.70 70.11 79.00 60.25 5.53
EvolveNav (Ours) 33.60 42.40 28.16 33.40 49.56 24.92 71.17 78.95 63.48 6.21
Unified Model for All Tasks:
NaviLLM [45] 44.56 53.74 36.63 35.44 - 28.09 67 - 58 5.91
NaviLLM* [45] 43.81 54.26 35.61 34.82 55.01 26.19 68.37 76.96 59.09 5.43
EvolveNav (Ours) 45.97 57.95 38.58 37.00 58.20 28.18 68.07 81.68 58.30 6.35
Cross-task training. Table V shows the cross-task training results on R2R [33], CVDN [34], SOON [36], and REVERIE [35]. From Table V, we can observe that EvolveNav surpasses the baseline approach NaviLLM [45] in most metrics on different benchmarks. These results reveal that our self-improving embodied reasoning framework is also effective for training the generalist navigation model, which is more practical and flexible in real-world navigation scenarios. Both the task-specific and cross-task training results on various VLN benchmarks sufficiently demonstrate that the proposed self-improving embodied reasoning framework fulfills adaptable and generalizable navigational reasoning under different tasks and scenarios.